Estimate the sample size needed to estimate a mean to within $135 with 99%
confidence if the standard deviation is $579.

123

71

62

0

244
Answer #1

Solution :

Given that,

Population standard deviation = = 579

Margin of error = E = 135

At 99% confidence level

= 1 - 99%  

= 1 - 0.99 =0.01

/2 = 0.005

Z/2 = Z0.005 = 2.576

  sample size = n = [Z/2* / E] 2

n = [ 2.576 * 579 / 135]2

n = 122.06

Sample size = n = 123